Operations Report – April 9th

Crew 314 Operations Report 09Apr2025

SOL: 3

Name of person filing report: Antoine Dubois (Crew Engineer)

Non-nominal systems :

Toilet was initially repaired at 11:50 a.m.

The water pipe disconnected again at 1:00 p.m.

Water supply was reconnected to the toilet at 4:30 p.m., and it seems to be holding.

Handbrake problem with the “Perseverance” rover.

Spacesuit #11 : A small part of the frigolite on space suit #11 was slightly damaged but has been repaired.

The suspected contact failure with headset #2 has been confirmed, 08Apr2025 around 9:30 p.m.

The CO detector in the science dome is not working.

Notes on non-nominal systems :

Toilet : All systems are functioning normally at this time.

After returning from the EVA on 09 Apr 2025, we noticed that the handbrake on rover “Perseverance” may no longer be holding properly.

Spacesuit #11 : During our EVA this morning, we noticed that a small piece of frigolite was damaged, but it was repaired when we arrived.

The headset with the contact defect has been replaced by a new one.

ROVERS

Spirit rover used: /

Hours: :/

Beginning charge: /

Ending charge: /

Currently charging: /

Opportunity rover used: EVA#5

Hours: 214.4 (0.8 used)

Beginning charge: 100%

Ending charge: 40 %

Currently charging: no

Curiosity rover used: /

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: /

Perseverance rover used: EVA #4

Hours: 318.1 H (0.8 used)

Beginning charge: 100%

Ending charge: 67%

Currently charging: no

General notes on rovers : Perserverance’s handbrake may no longer be holding properly. It was noticed at the end of the EVA, when coming back to the MDRS.

Summary of Hab operations: Today in the Habitat, we focused on cooking, socializing, completing our individual reports, and working on our respective experiments. Additionally, we carried out the second medical simulation as planned by Odile’s experiment (on the lower deck). Finally, we held our usual team meeting to discuss progress and upcoming tasks.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7:00 p.m. (Collection time has been brought forward by one hour for greater convenience)
1) Per formula (see notes): 27.58 gallons
2) Smart Home Dashboard (see notes): 29.01 gallons

Water (static tank, remaining gallons): 390.55 gallons (per formula) & 407.96 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : on

Toilet tank emptied (no or yes): yes

Summary of internet: used for music playing, mails, research for experiments, learning languages.

Summary of suits and radios:

Suits:

EVA# 5: 7, 8, 9, 10

Radio:

EVA # 5: 2,3,4, 7

Summary of GreenHab operations:

I watered all the plants with water. I took 30 garden labels for the experiment with pots and plants inoculated with arbuscular fungi.

The door of the GreenHab was left open for 3 hours 45 minutes (2.30 pm to 6.15 pm) to reduce the temperature inside (maximum of 97.5 °F)

WATER USE: 10.5 gallons

Heater (On or Off): Heater on (setup temperature = 72°F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 3 g of basilic

Summary of ScienceDome operations: put notes here:

Inoculation of Bacillus Thuringiensis on petri dish with soil or not and exposure to UV-C lamps

Inoculation of Esherichia Coli in 96 well plate and exposure to UV-C lamps

Preparation of material (pots with soil) to start the experiment of plant association with arbuscular fungi (after exposure to different stresses)

Check UV lamp operation and RPM (machine simulating the gravity)

Dual split (Heat or AC, On or Off): OFF

Summary of RAM operations: (Tools used, work done) put notes here

Crew Engineer took 2 colsons and a blue ribbon, at 8:50 a.m., to repair the equipment in the field if necessary, but nothing was broken. Everything has been put back in its place.

Crew Engineer and Executive Officer took an adjustable wrench, at 11:50 a.m., to fix the water leak in the toilet. Tools have been returned to their rightful place.

Executive Officer took a screwdriver, at 12:10 a.m., to fix the bike handlebars. Tools have been returned to their rightful place. It wasn’t properly attached to the exercise bike’s frame.

Crew Engineer and Executive Officer took an adjustable wrench, at 04:30 p.m., to reconnect the water supply to the toilet and it seems to be holding. Tools have been returned to their rightful place.

Summary of any observatory issues: no activities for the observatory

Summary of health and safety issues : Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support : put notes here

Crew requested the printing of labels crucial for one of their experiments

Supplemental Operations Report – March 29th

Date: 3/29/2025
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, invertors, generator. Robotic observatory. HAB outer shell, tunnels.

Power system:
Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and do not restart, which limits us to 5kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, Moreover, generattor is comsuming oil because of worn piston rings. Adding oil every day is necessary. Generator is limited to 8kW, see previous reports for details.
Main generator:
1) Oil, oil filter changed on 3/27/2025. Hours: 9045.6
2) Current hours – 9089.2

Propane Readings:
Did not notice refill
Station Tank: 79%
Director Tank: 77%
Intern Tank: 86%
Generator Tank: 56%

Water:
Hab Static Tank – 550 gallons
GreenHab – 200 gallons
Outpost tank – 550 gallons

Rovers:
Sojourner rover used: Yes
Hours: 212.5
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: None.

Cars:
Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: No.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should).

Summary of Internet: Nominal

EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Astro8 is malfunctioning and needs to be replaced.

Campus wide inspection, if action taken, what and why: Tunnel tarp repaired and will hopefully last until the end of the season.
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion.
Summary of GreenHab Operations: Nominal
Summary of SciDome Operations: Nominal
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Director’s Trailer furnace is malfunctioning intermittently. My guess is bad sensor(s) or problems with the logic board.
Summary of Health and Safety Issues: Nominal

HSO Pre-Mission Checklist – April 6th

HSO BEGINNING OF MISSION CHECKLIST 2024-2025

Submitted by: Odile Hilgers

Crew: 314

Date: April 6th 2025

Part 1

Locate and confirm the emergency escape routes in the Hab are functional and clear:

  1. Stairs (between lower end upper deck) : Clear and functional

  2. Emergency window (upper deck, east side) : Clear and functional

  3. Commander’s window (located in the commander’s crew quarter) : Clear an functional

Part 2

Inventory First Aid kit and note what needs to be refilled:

Main First Aid Cabinet (medical supplies and a few basic medications)

  • Blood Pressure Monitor (batteries OK)

  • Cotton Swabs (2 boxes)

  • Medical Alcohol 91% (⅓ of the first bottle, ⅔ of the second)

  • Antiseptic – Hydrogen Peroxide (2 bottles, nearly full)

  • Dental Floss (1 nearly full box)

  • Band-Aids (almost empty, ~10%) 🔁 [Needs refilling]

  • Sanitary Napkins (5)

  • Ibuprofen 200 mg (20 pills remaining out of 50)

  • Cutaneous Thermometer (batteries OK)

  • FFP2 Masks (12)

  • Standard Surgical Masks (13)

  • Emergency Blankets (4)

  • Emergency Splint (1)

  • Medical Tape (3 rolls, not full)

  • Various types of Bandages (7 rolls)

  • Motion Sickness Medication (1 unit)

  • Nitrile Gloves (1 pair)

  • Instant Cold Compress (unused / full)

Supplemental Operations Report -March 22nd

Date: 3/22/2025
Name of person filing report: David Steinhour
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, invertors, generator. Robotic observatory. HAB outer shell, tunnels.

Power system: "Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and do not restart, which limits us to 5kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, Moreover, generator is consuming oil because of worn piston rings. Adding oil every day is necessary. Generator is limited to 8kW, see previous reports for details.
Main generator:
1) Oil, oil filter, air filter changed on 3/19/2025.
2) Current hours – 8973.7"

Propane Readings: "Refilled 3/19/2025
Station Tank: 80%
Director Tank: 83%
Intern Tank: 86%
Generator Tank: 81%"

Water: "Hab Static Tank – 550 gallons
GreenHab – 200 gallons
Outpost tank – 550 gallons"

Rovers: "Sojourner rover used: Yes
Hours: 210.7
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: None."

Cars: "Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: No.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should)."

Summary of Internet: Guest wifi had to be turned on manually 3/20, but automation worked fine last night 3/21
EVA suits and radios: "Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Astro8 is malfunctioning and need to be replaced."
Campus wide inspection, if action taken, what and why: Tunnel tarp is once again ripped between science dome and solar observatory.
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion.
Summary of GreenHab Operations: Nominal
Summary of SciDome Operations: Nominal
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Nominal
Summary of Health and Safety Issues: Nominal

Supplemental Operations Report – March 16th

Date: 3/16/2025
Name of person filing report: David Steinhour
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, invertors, generator. Robotic observatory. HAB outer shell, tunnels.

Power system: "Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and did not restart, which limits us to 5kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, need to add oil every 2 days. Generator is limited to 8kW, see previous reports for details.
Main generator:
1) Oil, oil filter changed on 3/12/2025.
2) Current hours – 8874.0"

Propane Readings: "Refilled 2/26/2025
Station Tank: 57%
Director Tank: 49%
Intern Tank: 77%
Generator Tank: 34%"

Water: "Hab Static Tank – 550 gallons
GreenHab – 200 gallons
Outpost tank – 550 gallons"

Rovers: "Sojourner rover used: No
Hours: 209.5
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance tire patched 3/13, holding air well. Parking brake tightened. Battery maintenance performed on Sojourner and Perseverance."

Cars: "Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: No.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should)."

Summary of Internet: Nominal
EVA suits and radios: "Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Astro8 is malfunctioning and needs to be replaced."
Campus wide inspection, if action taken, what and why: Tunnel tarp is once again ripped between science dome and solar observatory.
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion.
Summary of GreenHab Operations: Nominal
Summary of SciDome Operations: Nominal
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Nominal
Summary of Health and Safety Issues: Nominal

Crew 309 HSO Beginning Of Mission Checklist

[title HSO Beginning Of Mission Checklist- March 17th]

Submitted by: Priyanka Vasu, Medic/Health and Safety Officer
Crew: 312
Date: 17Mar2025

Part 1
Locate and confirm the emergency escape routes in the Hab are functional and clear:
Stairs (between lower end upper deck): Functional and clear
Emergency window (upper deck, east side): Functional and clear
Commander’s window (located in the commander’s crew quarter): Functional and clear
Part 2
Inventory First Aid kit and note what needs to be refilled: Refilling is not required at the moment
Part 3
Note any safety issues: Nil
Note any health/environmental issues: Nil
Note any missing or recommended health and safety supplies: Nil
Part 4
Safety Equipment Inventory: All equipment are available and functional

Operations Report – March 17th

SOL: 1

Name of person filing report: Jesús, Engineer

Non-nominal systems: All systems nominal

Notes on non-nominal systems: No notes

ROVERS

Spirit rover : used

Beginning Hours: 268.8

Beginning charge: 73%

Ending Hours: 268.9

Ending charge: 72%

Currently charging: yes

Opportunity rover: used

Beginning Hours: 208.6

Beginning charge: 100%

Ending Hours: 209

Ending charge: 84%

Currently charging: no

Curiosity rover: Used

Beginning Hours: 301

Beginning charge: 95%

Ending Hours: 301.7

Ending charge: 71%

Currently charging: yes

Perseverance rover:

Beginning Hours: 311

Beginning charge: 93%

Ending Hours: 311.4

Ending charge: 77%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Equipment troubleshooting, exercise activities, cooking

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 4:50 pm
1) Per formula (see notes): 111.1 gal
2) Smart Home Dashboard (see notes): 71.63 gal

Water (static tank, remaining gallons): 438.91 gal

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: used from 7 pm to 10 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, used

Suits : For EVA 2, had to use MDRS suit due to equipment malfunction.

Radios : Nothing to report.

Summary of GreenHab operations: Plants were watered.

WATER USE: 5 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams):

Basil: 4 g

Cherry tomatoes: 80 g

Radishes: 115 g

Scallions: 57 g

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): None

Summary of any observatory issues: None

Summary of health and safety issues: None

Operations Report – March 18th

SOL: 2

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: HAB Heating system

Notes on non-nominal systems: The HAB heating system was failing to heat consistently due to excessive winds, replaced air filter, checked for gas leaks, replaced air filter, shut off heating system at around 3 pm due to intense propane smell across the HAB and began use of the electric space heater, but returned to furnace use around 6 pm.

ROVERS

Spirit rover : not used

Opportunity rover: not used

Curiosity rover: Used

Beginning Hours: 301.7

Beginning charge: 100%

Ending Hours: 302

Ending charge: 73%

Currently charging: no

Perseverance rover: used

Beginning Hours: 311.5

Beginning charge: 100%

Ending Hours: 311.8

Ending charge: 75%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Heating system troubleshooting, cooking, EVA preparation

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 4:50 pm
1) Per formula (see notes): 45.75 gal
2) Smart Home Dashboard (see notes): 27.35 gal

Water (static tank, remaining gallons): 451.02 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: used from 7 pm to 10 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Nothing to report.

Radios : During EVA 5, chatter between radio 4 and 5 was full of static for around 30 min of the EVA

Summary of GreenHab operations: Plants were watered.

WATER USE: 12 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams):

Basil: 30 g

Carrots: 92 g

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): Medic worked on medical rover, no specific tools were used. Gas sniffer and space heater were moved from RAM to the main HAB due to heating issues.

Summary of any observatory issues: None

Summary of health and safety issues: None

Operations Report – March 19th

SOL: 3

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover : not used

Opportunity rover: not used

Curiosity rover: Used

Beginning Hours: 302

Beginning charge: 100%

Ending Hours: 302.4

Ending charge: 75%

Currently charging: yes

Perseverance rover: used

Beginning Hours: 311.8

Beginning charge: 100%

Ending Hours: 312

Ending charge: 92%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Exercise activities, cooking, EVA preparation, equipment troubleshooting

Water Use (please use both methods to estimate water usage)

Time of measurements: 5:00 pm
1) Per formula (see notes): 22.22 gal
2) Smart Home Dashboard (see notes): 32.83 gal

Water (static tank, remaining gallons): 418.19 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: used from 7 pm to 10 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Bolts on helmet hinges were very loose during/after the first EVA but were tightened for the second.

Radios : Radio 1 stopped working during the last portion of second EVA but resumed functionality after it was turned off and on upon return to HAB. Cause unknown.

Summary of GreenHab operations: Plants were watered.

WATER USE: 6 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams): None

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): None

Summary of any observatory issues: None

Summary of health and safety issues: None

Operations Report – March 20th

SOL: 4

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: Wi-fi

Notes on non-nominal systems: Wi-fi failed to turn on during the regular window, believe the cause was due to the outages from earlier in the day.

ROVERS

Spirit rover : not used

Opportunity rover: not used

Curiosity rover: used

Beginning Hours: 302

Beginning charge: 100%

Ending Hours: 302.5

Ending charge: 88%

Currently charging: yes

Perseverance rover: used

Beginning Hours: 312

Beginning charge: 100%

Ending Hours: 312.3

Ending charge: 78%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Exercise activities, cooking, EVA preparation, equipment troubleshooting

Water Use (please use both methods to estimate water usage)

Time of measurements: 5:00 pm
1) Per formula (see notes): 28.75 gal
2) Smart Home Dashboard (see notes): 42.27 gal

Water (static tank, remaining gallons): 375.92 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): yes

Summary of internet: will be using from 9 pm to 11 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Nothing to report.

Radios : Nothing to report.

Summary of GreenHab operations: Plants were watered.

WATER USE: 7 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams): None

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): None

Summary of any observatory issues: None

Summary of health and safety issues: None

Copyright © The Mars Society. All rights reserved. | Main Site